OpenStudy (anonymous):

x − 3y if x = −3 and y = −2

OpenStudy (oaktree):

-3 - 3(-2) = -3+6=3

OpenStudy (anonymous):

Huh?

OpenStudy (diyadiya):

Plug in the values of x & y in the equation x-3y x=-3 y=-2 (-3)-3(-2)

OpenStudy (anonymous):

what would it equal

OpenStudy (diyadiya):

-3-3(-2) = -3 - (-6) = ? what do you get?

OpenStudy (anonymous):

Use the distributive property to simplify the expression below: −4(3u − 5)